# Neural Spacetimes for DAG Representation Learning

### v1

@@ 2.2 Spatial Structure
Quasi-Metric Spaces. Although Riemannian manifolds have been employed to formalize non-Euclidean distances between points, their additional structural characteristics, such as smoothness and infinitesimal angles, present considerable constraints. This complexity often makes it challenging to express the distance function for arbitrary Riemannian manifolds in closed-form. On the other hand, quasi-metric spaces isolate the pertinent properties of Riemannian distance functions without requiring any of their additional structures for graph embedding. A quasi-metric space is defined as a set $X$ equipped with a distance function $d:X\times X\rightarrow[0,\infty)$ satisfying the following conditions for every $x_{u},x_{v},x_{w}\in X$: i) $d(x_{u},x_{v})=0$ if and only if $x_{u}=x_{v}$, ii) $d(x_{u},x_{v})=d(x_{v},x_{u})$, iii) $d(x_{u},x_{v})\leq C\big{(}d(x_{u},x_{w})+d(x_{w},x_{v})\big{)}$, for some constant $C\geq 1$. Property (iii) is called the $C$-relaxed triangle inequality. When $C=1$, $(X,d)$ is termed a metric space, and the $C$-relaxed triangle inequality becomes the standard triangle inequality. Quasi-metrics arise naturally when considering local embeddings since the triangle inequality is only required to hold locally, allowing for small neighbourhoods of distinct points in a point metric space to be embedded independently from one another. Furthermore, these geometries share many of the important properties of metric spaces, e.g. the Arzela-Ascoli theorem holds (Xia, 2009). Moreover, if property (i) is dropped, such that several points may be indistinguishable by their distance information, then we say that $d$ is a pseudo-quasi-metric (Kim, 1968; Kũnzi, 1992). This naturally occurs when additional time dimensions are used to encode causality, but distance is ignored.
@@
$d_{G}(u,v)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf\Biggl{\{}\sum_{n=1}^{N-1}\,W(\nu_{n},\nu_{n+1}):\exists\,(\nu_{1}=u,\nu_{2}),\dots,(\nu_{N-1},\nu_{N}=v)\in E\Biggr{\}}.$ (1)
@@ 2.3 Spacetime Embeddings
(ii) Metric Embedding: $c\,\rho(x_{1},x_{2})\leq d(x_{1},x_{2})\leq D\,c\rho(x_{1},x_{2}).$ The constant $D$ is called the distortion, and $c$ is the scale of the spacetime embedding. If $D=c=1$, then $f$ is called isocausal.
Our description of spacetime embeddings is illustrated in Figure 3. We encode causality using the notion of an order embedding from order theory (Davey & Priestley, 2002), also used in general relativity (Kronheimer & Penrose, 1967; Sorkin, 1991; Reid, 2003; Henson, 2009; Benincasa & Dowker, 2010), jointly with the notion of a metric embedding core to theoretical computation science (Gupta & Hambrusch, 1992; Gupta, 1999; Gupta et al., 2004) and representation learning (Sonthalia & Gilbert, 2020).
@@ 3 Neural Spacetimes
Problem Formulation.
Following the discussion in Section 2, our goal is to learn a spacetime manifold representation where nodes in a DAG can be embedded as events. This should be done while capturing edge directionality in the form of causal structure (which includes modelling lack of connectivity as anti-chains), as well as the edge weights in the original discrete structure: these are represented as temporal ordering and spatial distance in the continuous embedding space.
3.1 The Neural Spacetime Architecture
Let $N\in\mathbb{N}_{+}$ be the dimensionality of the feature vectors $x_{u},x_{v}\in\mathbb{R}^{N}$ associated with nodes $u,v\in V$, where $V$ represents the set of nodes of the DAG $G_{D}=(E_{D},V,W_{D})$. Fix a space dimension $D\in\mathbb{N}_{+}$ and a time dimension $T\in\mathbb{N}_{+}$. A neural spacetime (NST) is a learnable triplet $\mathcal{S}=(\mathcal{E},\mathcal{D},\mathcal{T})$, where $\mathcal{E}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{D+T}$ is an encoder network, e.g. an MLP, $\mathcal{D}:\mathbb{R}^{D+T}\times\mathbb{R}^{D+T}\to[0,\infty)$ is a learnable quasi-metric on $\mathbb{R}^{D}$ and $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ is a learnable partial order on $\mathbb{R}^{T}$. We implement $\mathcal{D}$ and $\mathcal{T}$ as two neural networks that process the spatial and temporal dimensions of encoded feature vectors $\hat{x}_{u},\hat{x}_{v}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\in\mathbb{R}^{D+T}$ in parallel.
@@ 3.1 The Neural Spacetime Architecture
To define $\mathcal{D}$, which will be implemented using a variation of the original neural snowflake (see equation 10 from Appendix A.3), we consider the following activation function, $\sigma_{s,l}(x)$. The activation depends on two trainable parameters $s,l>0$. In the spirit of the fractalesque structure of neural snowflakes, $s$ controls whether small-scale distances should be expanded or contracted relative to the distance on $\mathbb{R}^{D}$. Similarly, $l$ controls the large-scale distances of points and dictates whether those should be expanded or contracted. See Appendix C.1 for an extended discussion.
@@
$\sigma_{s,l}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\begin{cases}\operatorname{sgn}(x)\,|x|^{s}&\mbox{ if }|x|<1\\ \operatorname{sgn}(x)\,|x|^{l}&\mbox{ if }|x|\geq 1\end{cases}$ (2)
@@
$\mathcal{D}(\hat{x}_{u},\hat{x}_{v})$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathsf{W}_{J}\sigma_{s_{J},l_{J}}\bullet(u_{J-1});\,u_{j}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathsf{W}_{j}\sigma_{s_{j},l_{j}}\bullet(u_{j-1})\mbox{ for }j=1,\dots,J-1,$ (3)
$u_{0}$ =$\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}|\sigma_{s_{0},l_{0}}\bullet(\hat{x}_{u})_{1:D}-\sigma_{s_{0},l_{0}}\bullet(\hat{x}_{v})_{1:D}|,$
for each $\hat{x}_{u},\hat{x}_{v}\in\mathbb{R}^{D+T}$, where the depth parameter $J\in\mathbb{N}_{+}$, where the $s_{0},l_{0},\dots,s_{J},l_{J}>0$, and $\bullet$ denotes component-wise composition, weight matrices $\mathsf{W}_{j}\in I^{+}_{D}$ (an invertible positive matrix, see Appendix A.2) for $j<J$ and $\mathsf{W}_{J}\in(0,\infty)^{1\times d}$ all of which have positive entries, and where the absolute value ||$\cdot$|| ⋅ | is also applied component-wise. Note that $(\cdot)_{1:D}$ extracts the first $D$ dimensions of the input feature vector.
Moreover, we seek an ordering $\lesssim$ on $\mathbb{R}^{D+T}$ so that $u\preccurlyeq v$ in the poset if and only if their respective embeddings $\hat{x}_{u}$ and $\hat{x}_{v}$ are ordered in the feature space via $\lesssim$. Our class of trainable orders are parameterized with the following type of neural networks, which we call neural partial orders. Consider a map $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ admitting the iterative representation
$\mathcal{T}(\hat{x}_{u})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}z_{\tilde{J}},~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall j\in\{1,\dots,\tilde{J}\},z_{j}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathsf{V}_{j}\sigma_{\tilde{s}_{j},\tilde{s}_{j}}\circ\textrm{LR}\bullet(z_{j-1})+b_{j},~{}~{}~{}~{}~{}~{}~{}~{}~{}z_{0}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\hat{x}_{u})_{D+1:D+T},$ (4)
@@
If $T\in\mathbb{N}_{+}$, $D=0$, and $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ admits a representation as equation 4, then $\lesssim^{\mathcal{T}}$ is a partial order on $\mathbb{R}^{D+T}$.
Proof.
See page $21$. ∎
@@
Any $\mathcal{D}$ with representation in equation 3 is a quasi-metric on $\mathbb{R}^{D}$. If each $\mathsf{W}_{j}$ is orthogonal  [^1] Since any $\mathsf{W}_{j}$ must have only non-negative entries, then if $\mathsf{W}_{j}$ is additionally orthogonal, it must be a permutation matrix. and $0<s_{j},t_{j}\leq 1$ then $\mathcal{D}$ is a metric on $\mathbb{R}^{D}$. Furthermore, if $T\in\mathbb{N}_{+}$, then $\mathcal{D}$ is a pseudo-quasi-metric on $\mathbb{R}^{D+T}$.
Proof.
See page $21$. ∎
@@ 3.2 Embedding Guarantees
Let $(P,\lesssim)$ be a poset. A subset $S\subseteq P$ is called an anti-chain if for each $u,v\in S$ neither $u\preccurlyeq v$ nor $v\preccurlyeq u$. The width of $(P,\lesssim)$ is the maximal size (cardinality) of any anti-chain in $P$.
@@
Fix $W,N\in\mathbb{N}_{+}$, $K>0$, and let $(P,d,\lesssim)$ be a finite causal metric space with doubling constant $K$, width $W$, and $P\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x_{v}\}_{v\in V}\subseteq\mathbb{R}^{N}$. There exists a $D\in\mathcal{O}(\log(K))$, an MLP $\mathcal{E}:\mathbb{R}^{N}\to\mathbb{R}^{D+W}$, $\mathcal{T}:\mathbb{R}^{D+W}\to\mathbb{R}^{W}$ with representation equation 4, and $\mathcal{D}:\mathbb{R}^{D+W}\times\mathbb{R}^{D+W}\to[0,\infty)$ with representation equation 3 such that for each $x_{u},x_{v}\in P$
@@
Moreover, setting $D=k$, (ii) can be improved to an isocausal embedding, i.e. $d(x_{u},x_{v})=\mathcal{D}\big{(}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\big{)}$. In either case, the number of non-zero parameters determining the neural spacetime triplet $\mathcal{S}=(\mathcal{E},\mathcal{D},\mathcal{T})$ is $\tilde{\mathcal{O}}\big{(}D+W+k^{5/2}D^{4}N\big{)}.$
Proof.
See page $25$. ∎
@@
If the Hasse diagram of a poset $(P,\lesssim)$ is not a planar graph, then there is no spacetime embedding $\mathcal{E}:P\to\mathbb{R}^{D+W}$, $\mathcal{D}:\mathbb{R}^{D+W}\times\mathbb{R}^{D+W}\to[0,\infty)$, $\mathcal{T}:\mathbb{R}^{D+W}\to\mathbb{R}^{W}$ with $W=1,2$.
Proof.
See page $26$. ∎
@@
(i) Causal Embedding: $x_{u}\preccurlyeq x_{v}$ if and only if $\mathcal{E}(x_{u})\lesssim^{\mathcal{T}}\mathcal{E}(x_{v})$,
(ii) Low-Distortion Metric Embedding: $d_{\operatorname{H}}(x_{u},x_{v})\leq\mathcal{D}\big{(}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\big{)}\leq\mathcal{O}(\log(k)^{2})\,d_{\operatorname{H}}(x_{u},x_{v}).$
The total number of parameters in $\mathcal{S}$ is $\mathcal{O}\big{(}k^{5/2}D^{4}N\,\log(N)\,\,\log\big{(}k^{2}\,\operatorname{diam}(P,d_{\operatorname{H}})\big{)}\big{)}.$
Proof.
See page $26$. ∎
@@
Remark 3 (Local vs. Global NST Embeddings.).
@@
Lastly, for completeness we also provide a discussion comparing neural spacetimes to hyperbolic representation in Appendix A.6.

### v2

@@ 2.2 Spatial Structure
Quasi-Metric Spaces. Although Riemannian manifolds have been employed to formalize non-Euclidean distances between points, their additional structural characteristics, such as smoothness and infinitesimal angles, present considerable constraints. This complexity often makes it challenging to express the distance function for arbitrary Riemannian manifolds in closed-form. On the other hand, quasi-metric spaces isolate the pertinent properties of Riemannian distance functions without requiring any of their additional structures for graph embedding. A quasi-metric space is defined as a set $X$ equipped with a distance function $d:X\times X\rightarrow[0,\infty)$ satisfying the following conditions for every $x_{u},x_{v},x_{w}\in X$: i) $d(x_{u},x_{v})=0$ if and only if $x_{u}=x_{v}$, ii) $d(x_{u},x_{v})=d(x_{v},x_{u})$, iii) $d(x_{u},x_{v})\leq C\big{(}d(x_{u},x_{w})+d(x_{w},x_{v})\big{)}$, for some constant $C\geq 1$. Property (iii) is called the $C$-relaxed triangle inequality. When $C=1$, $(X,d)$ is termed a metric space, and the $C$-relaxed triangle inequality becomes the standard triangle inequality. Quasi-metrics arise naturally when considering local embeddings since the triangle inequality is only required to hold locally, allowing for small neighborhoods of distinct points in a point metric space to be embedded independently from one another. Furthermore, these geometries share many of the important properties of metric spaces, e.g. the Arzela-Ascoli theorem holds (Xia, 2009). Moreover, if property (i) is removed, such that several points may be indistinguishable by their distance information, then we say that $d$ is a pseudo-quasi-metric (Kim, 1968; Kũnzi, 1992). This naturally occurs when additional time dimensions are used to encode causality, but distance is ignored.
@@
$d_{G}(u,v)\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\inf\Biggl{\{}\sum_{n=1}^{N-1}\,W(\nu_{n},\nu_{n+1}):\exists\,(\nu_{1}=u,\nu_{2}),\dots,(\nu_{N-1},\nu_{N}=v)\in E\Biggr{\}}.$ (1)
@@ 2.3 Spacetime Embeddings
(ii)Metric Embedding: $\color[rgb]{0,0,0}{c\,\rho(f(x_{1}),f(x_{2}))\leq d(x_{1},x_{2})\leq D\,c\rho(f(x_{1}),f(x_{2})).}$
The constant $D$ is called the distortion, and $c$ is the scale of the spacetime embedding. A spacetime embedding $f:\mathcal{X}\to\mathcal{Y}$ for which $D=c=1$ is called isocausal.
Remark 2 (Optimal Distortion).
Note that $D=1$ is the minimal possible distortion [^1] The symbol $D$ is also used in other parts of the text to denote the spatial dimensionality of the NST. We stick to $D$ here since it is standard in the metric embedding literature, but the two should not be confused.. This is because $D<1$ is impossible and $D=1$ yields an equality between the rescaled target metric $c\,\rho$ and the original metric $d$; i.e. $c\rho(f(x_{1}),f(x_{2}))=d(x_{1},x_{2})$ for all $x_{1},x_{2}\in\mathcal{X}$.
Our spacetime embeddings are illustrated in Figure 3. We encode causality using the notion of an order embedding from order theory (Davey & Priestley, 2002), also used in general relativity (Kronheimer & Penrose, 1967; Sorkin, 1991; Reid, 2003; Henson, 2009; Benincasa & Dowker, 2010), jointly with the notion of a metric embedding core to theoretical computation science (Gupta & Hambrusch, 1992; Gupta, 1999; Gupta et al., 2004) and representation learning (Sonthalia & Gilbert, 2020).
@@ 3 Neural Spacetimes
Let $N\in\mathbb{N}_{+}$ be the dimensionality of the feature vectors $x_{u},x_{v}\in\mathbb{R}^{N}$ associated with nodes $u,v\in V$, where $V$ represents the set of nodes of the DAG $G_{D}=(E_{D},V,W_{D})$. Fix a space dimension $D\in\mathbb{N}_{+}$ and a time dimension $T\in\mathbb{N}_{+}$. A neural spacetime (NST) is a learnable triplet $\mathcal{S}=(\mathcal{E},\mathcal{D},\mathcal{T})$, where $\mathcal{E}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{D+T}$ is an encoder network, e.g. an MLP, $\mathcal{D}:\mathbb{R}^{D+T}\times\mathbb{R}^{D+T}\to[0,\infty)$ is a learnable quasi-metric on $\mathbb{R}^{D}$ and $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ is a learnable partial order on $\mathbb{R}^{T}$. We implement $\mathcal{D}$ and $\mathcal{T}$ as two neural networks that process the spatial and temporal dimensions of encoded feature vectors $\hat{x}_{u},\hat{x}_{v}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\in\mathbb{R}^{D+T}$ in parallel.
@@
To define $\mathcal{D}$, which will be implemented using a variation of the original neural snowflake (see equation 10 from Appendix A.3). The original neural snowflake, and our upgrade thereof, employs two neural networks: the first represents the metric space as vectorial data in $\mathbb{R}^{D}$, and the second perturbs the Euclidean distance thereon. Together, any finite metric space may be perfectly (isometrically) embedded in this manner, whereas only one of the networks alone is not enough; e.g. expander graphs cannot be isometrically embedded into any low-dimensional Euclidean space (Kratsios et al., 2023a, Proposition 13).
We consider the following activation function, $\sigma_{s,l}$. The activation $\sigma_{s,l}$ depends on two trainable parameters $s,l>0$. In the spirit of the fractalesque structure of neural snowflakes, $s$ controls whether small-scale distances should be expanded or contracted relative to the distance on $\mathbb{R}^{D}$. Similarly, $l$ controls the large-scale distances of points and dictates whether those should be expanded or contracted. See Appendix C.1 for an extended discussion.
@@
$\sigma_{s,l}(x)\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\begin{cases}\operatorname{sgn}(x)\,|x|^{s}&\mbox{ if }|x|<1\\ \operatorname{sgn}(x)\,|x|^{l}&\mbox{ if }|x|\geq 1\end{cases}$ (2)
@@
$\mathcal{D}(\hat{x}_{u},\hat{x}_{v})$ =$\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\mathsf{W}_{J}\sigma_{s_{J},l_{J}}\bullet(u_{J-1});\,u_{j}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\mathsf{W}_{j}\sigma_{s_{j},l_{j}}\bullet(u_{j-1})\mbox{ for }j=1,\dots,J-1,$ (3)
$u_{0}$ =$\stackrel{{\scriptstyle\mbox{def.}}}{{=}}|\sigma_{s_{0},l_{0}}\bullet(\hat{x}_{u})_{1:D}-\sigma_{s_{0},l_{0}}\bullet(\hat{x}_{v})_{1:D}|,$
for each $\hat{x}_{u},\hat{x}_{v}\in\mathbb{R}^{D+T}$, where the depth parameter $J\in\mathbb{N}_{+}$, where the $s_{0},l_{0},\dots,s_{J},l_{J}>0$, and $\bullet$ denotes component-wise composition (i.e. the function is applied element-wise, which is standard in Deep Learning), weight matrices $\mathsf{W}_{j}\in I^{+}_{D}$ (an invertible positive matrix, see Appendix A.2 ) for $j<J$ and $\mathsf{W}_{J}\in(0,\infty)^{1\times d}$ all of which have positive entries, and where the absolute value ||$\cdot$|| ⋅ | is also applied component-wise. Note that $(\cdot)_{1:D}$ extracts the first $D$ dimensions of the input vector.
Moreover, we seek an ordering $\lesssim$ on $\mathbb{R}^{D+T}$ so that $u\preccurlyeq v$ in the poset if and only if their respective embeddings $\hat{x}_{u}$ and $\hat{x}_{v}$ are ordered in the feature space via $\lesssim$. Our class of trainable orders are parameterized with the following type of neural networks, which we call neural partial orders.
Consider a map $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ admitting the iterative representation
$\mathcal{T}(\hat{x}_{u})\stackrel{{\scriptstyle\mbox{def.}}}{{=}}z_{\tilde{J}},~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall j\in\{1,\dots,\tilde{J}\},z_{j}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\mathsf{V}_{j}\sigma_{\tilde{s}_{j},\tilde{s}_{j}}\circ\textrm{LR}\bullet(z_{j-1})+b_{j},~{}~{}~{}~{}~{}~{}~{}~{}~{}z_{0}\stackrel{{\scriptstyle\mbox{def.}}}{{=}}(\hat{x}_{u})_{D+1:D+T},$ (4)
@@
If $T\in\mathbb{N}_{+}$, $D=0$, and $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ admits a representation as equation 4, then $\lesssim^{\mathcal{T}}$ is a partial order on $\mathbb{R}^{D+T}$. See page $21$ for proof.
@@
Any $\mathcal{D}$ with representation in equation 3 is a quasi-metric on $\mathbb{R}^{D}$. If each $\mathsf{W}_{j}$ is orthogonal  [^2] Since any $\mathsf{W}_{j}$ must have only non-negative entries, then if $\mathsf{W}_{j}$ is additionally orthogonal, it must be a permutation matrix. and $0<s_{j},l_{j}\leq 1$ then $\mathcal{D}$ is a metric on $\mathbb{R}^{D}$.
Furthermore, if $T\in\mathbb{N}_{+}$, then $\mathcal{D}$ is a pseudo-quasi-metric on $\mathbb{R}^{D+T}$. See page $21$ for proof.
@@ 3.1 Embedding Guarantees
Let $(P,\lesssim)$ be a poset. A $S\subseteq P$ is called an anti-chain if for each $u,v\in S$ neither $u\preccurlyeq v$ nor $v\preccurlyeq u$. The width of $(P,\lesssim)$ is the maximal cardinality of any anti-chain in $P$.
@@
Fix $W,N\in\mathbb{N}_{+}$, $K>0$, and let $(P,d,\lesssim)$ be a finite causal metric space with doubling constant $K$, width $W$, and $P\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\{x_{v}\}_{v\in V}\subseteq\mathbb{R}^{N}$. There exists a $D\in\mathcal{O}(\log(K))$, an MLP $\mathcal{E}:\mathbb{R}^{N}\to\mathbb{R}^{D+W}$, $\mathcal{T}:\mathbb{R}^{D+W}\to\mathbb{R}^{W}$ with representation equation 4, and $\mathcal{D}:\mathbb{R}^{D+W}\times\mathbb{R}^{D+W}\to[0,\infty)$ with representation equation 3 such that for each $x_{u},x_{v}\in P$
@@
Moreover, setting $D=k$, (ii) can be improved to an isocausal embedding, i.e. $d(x_{u},x_{v})=\mathcal{D}\big{(}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\big{)}$. In either case, the number of non-zero parameters determining the neural spacetime triplet $\mathcal{S}=(\mathcal{E},\mathcal{D},\mathcal{T})$ is $\tilde{\mathcal{O}}\big{(}D+W+k^{5/2}D^{4}N\big{)}.$ See page $25$ for proof.
@@
If the Hasse diagram of a poset $(P,\lesssim)$ is not a planar graph, then there is no spacetime embedding $\mathcal{E}:P\to\mathbb{R}^{D+W}$, $\mathcal{D}:\mathbb{R}^{D+W}\times\mathbb{R}^{D+W}\to[0,\infty)$, $\mathcal{T}:\mathbb{R}^{D+W}\to\mathbb{R}^{W}$ with $W=1,2$. See page $26$ for proof.
@@
(i) Causality: $x_{u}\preccurlyeq x_{v}$ if and only if $\mathcal{E}(x_{u})\lesssim^{\mathcal{T}}\mathcal{E}(x_{v})$,
(ii) Low-Distortion: $d_{\operatorname{H}}(x_{u},x_{v})\leq\mathcal{D}\big{(}\mathcal{E}(x_{u}),\mathcal{E}(x_{v})\big{)}\leq\mathcal{O}(\log(k)^{2})\,d_{\operatorname{H}}(x_{u},x_{v}).$
The number of parameters in $\mathcal{S}$ is $\mathcal{O}\big{(}k^{5/2}D^{4}N\,\log(N)\,\,\log\big{(}k^{2}\,\operatorname{diam}(P,d_{\operatorname{H}})\big{)}\big{)}.$ Proof on page $26$.
@@
Remark 4 (Local vs. Global NST.).
@@
For completeness, we compare neural spacetimes to hyperbolic representation in Appendix A.6.